# Elasto-frictional reduced model of a cyclically sheared container filled with particles

### v1

@@ 4.1 Effective elasticity
It is observed in Fig. 4a that the identified elastic moduli $G_{D}$ and $G_{E}$ are both correlated to the confinement pressure $P$. Such a behavior is expected in granular material. Indeed, focusing on the low amplitude regimes, such that $\tau_{D}\ll\mu_{D}P$, the asymptote of the reduced model defined in Eq. 4 and 5 indicates that ${\tau\rightarrow(G_{E}+G_{D})\gamma}$. Hence, a total shear modulus of the granular packing,
@@
is enhanced due to the contribution of the inter-particles contact stiffness in the tangential direction, $k_{t}$, defined in Eq. 2. The expressions Eqs. 10-12 indicates the typical scaling $G_{\mathrm{EMT}}^{0,\infty}\propto K_{\mathrm{EMT}}\propto E_{p}(P/E_{p})^{1/3}$ inherited Hertzian contact interaction. Finally, it is worth reminding that the volume fraction and the coordination number are also pressure dependent (O’Hern et al., 2003), $\phi(P)$ and $Z(P)$, as pointed out in Sec. 3.5.
@@
with a well-defined coefficient of proportionality, $\alpha=0.687\pm 0.073~{}(10.6\%)$, in average over a dataset which is restricted to an inter-particles friction coefficient $\mu_{p}=0.15$. In line with the EMT model, the value of the prefactor relies on the inter-particle friction coefficient: the formula given by Eqs. 11-12 indicate two asymptotic values $\alpha=G_{\mathrm{EMT}}^{0}/G_{\mathrm{EMT}}^{\infty}=1/(1+3k_{t}/2k_{n})\simeq 0.4$ for frictionless particles and $\alpha=G_{\mathrm{EMT}}^{\infty}/G_{\mathrm{EMT}}^{\infty}=1$ for frictional ones. The value of $\alpha$ at an intermediate $\mu_{p}=0.15$ is thus consistent with these bounds.
Figure 5: (a): Total shear modulus $G_{T}=G_{D}+G_{E}$ versus effective shear modulus $G_{\mathrm{EMT}}^{\infty}$ (Eq. 12) with a linear fit of slope $\alpha=0.687\pm 0.073~{}(10.6\%)$. Same legend as in Fig. 4. (b,c): identified Dahl friction coefficient $\mu_{D}$ and coefficient $\alpha$ as functions of the inter-particles friction $\mu_{p}$.
@@ 4.2 Effective friction
which is known to depend on a single dimensionless inertial number (Da Cruz et al., 2005; Jop et al., 2006), $I=t_{p}/t_{s}$. The number $I$ is the ratio between two time scales, namely, the typical time $t_{s}=1/\dot{\gamma}$ associated to shear rate, which correspond to the macroscopic deformation duration, and $t_{p}=r_{p}/\sqrt{P/\rho_{p}}$ the one related to confinement, corresponding to the typical duration of microscopic rearrangement during which a grain moves by one neighbor size due the confining pressure (Jop, 2015). The ratio $I$ describes the regime of a granular flow, from the quasi-static limit at low $I$ to the kinetic regime with multiple collisions at high $I$. A phenomenological equation (Jop et al., 2006) provides an estimation of the effective friction coefficient,
@@
where $I_{0}$ is a constant of the order of unity, which delimits the two asymptotic limits. The effective friction coefficient value $\mu_{g}$ accounts for the existence of a gaseous phase at high inertia (Jop, 2015), $I\gg I_{0}$. In turn, the quasi-static value of the effective friction coefficient, $\mu_{s}$, mostly depends on the inter-particle friction coefficient $\mu_{p}$ (Badetti et al., 2018). In practice, the estimated inertial number probed in our study remains small, $10^{-5}\leq I\leq 10^{-4}$, thanks to the low driving frequency (see Sec. 2), in such a way that the effective friction does not depend on the inertial number but only on the inter-particle friction coefficient, $\mu(I\ll I_{0})\simeq\mu_{s}(\mu_{p})$. Consistently, the analysis of the dependence of the Dahl’s effective friction coefficient, $\mu_{D}$, as a function of the inter-particle one, $\mu_{p}$, is shown on Fig. 5b, all other parameters being fixed. In particular, we observe that $\mu_{D}$ is an increasing function of $\mu_{p}$ with a saturation, in qualitative and quantitative agreements with the literature, see for instance Ref. (Badetti et al., 2018).
Finally, we check the dependence of the prefactor $\alpha$ defined in Eq. 13 with respect to $\mu_{p}$. As expected in Sec. 4.1, $\alpha$ should be a function of $\mu_{p}$, accounting for the transition between perfectly frictionless, $\alpha\simeq 0.4$, and infinitely frictional particles, $\alpha=1$. The observation reported in Fig. 5c is in qualitative agreement with this expectation, showing that the prefactor $\alpha$ increases with the inter-particle friction coefficient $\mu_{p}$ until saturating to a value of the order of unity.
@@ 4.3 Dimensional analysis
We determined in Sec. 3.5 that the mechanical response depends on numerous particles features ($E_{p}$, $\nu_{p}$, $\rho_{p}$, $r_{p}$, $N_{p}$ and $\mu_{p}$), state variables ($P$, $Z$, $\phi$) and external parameters ($\omega_{0}$ and $\gamma_{m}$). However, Sec. 4.1 and 4.2 show that most of these parameters are linked with each other. Here, we aim at reducing the dimension of the parameters space, by making use of the observations presented in Sec. 3.3. In particular, we use the fact that the stress strain response is well described by an elasto-hysteretic model given by Eq. 4 and 5. We thus choose to rewrite these equations in dimensionless form using the following quantities:
@@
with $\Gamma$ and $\mathrm{T}$ the dimensionless shear strain and stress, respectively normalized by the amplitude of the shear strain and a typical shear stress proportional to the total shear modulus $G_{T}$ defined in Eq. 9. In line with this choice, it stands to reason that the Dahl’s effective friction coefficient $\mu_{D}$ can be represented by a non-dimensional parameter $\xi$, the first parameter of our model, resulting from the competition between the typical shear stress imposed to the granular medium, $G_{T}\gamma_{m}$, and a Mohr-Coulomb-like threshold $\mu_{D}P$. According to Fig. 4, the parameter $\xi$ is essentially a function of $P/k_{n}$ and $\gamma_{m}$. In turn, the elastic term added to the hysteretic contribution, see Eqs. 4 and 5, can be described by a second parameter, $\Psi$, which stands as a partition parameter quantifying the ratio between the purely elastic shear modulus, $G_{E}$, and the one stemming from the hysteretic response, $G_{D}$. According to the normalizations defined in Eq. 16, the model Eqs. 4 and 5 rewrites
@@
revealing that the mechanical response $\mathrm{T}(\Gamma)$ only depends on two parameters, $\xi$ and $\Psi$. From a physical point of view, the parameter $\xi$ provides an information on the degree of hysteresis (related to frictional stress, $\mu_{D}P$) and elasticity (related to elastic stress, $G_{T}\gamma_{m}$) of the response, suggesting that $\Psi$ should be also a function of $\xi$.
@@
Such an expectation is checked in Fig. 6, which reveals a correlation between $\Psi$ and $\xi$. In practice, the dependency is probed in the small-to-moderate values of $\xi$ only (see Fig. 1d for instance), where an hysteresis is clearly observable and leads to accurate estimations of the triplet $(\mu_{D},G_{D},G_{E})$ i.e. of a pair $(\Psi,\xi)$. For larger $\xi$ (i.e. high pressure $P$ or low amplitude $\gamma_{m}$), the mechanical response tends to an elastic one, which prevents distinguishing the elasto-hysteretic contributions and leads to poorly accurate estimations of $\Psi$ and $\xi$. However, considering that the granular medium becomes purely elastic in the high $\xi$ regime provides the information that the partition parameter $\Psi$ saturates to 1. Hence, the simplest saturating function takes the form
@@
where $\beta$ is an adjusting parameter. Fitting the Ansatz given in Eq. 18 to both $G_{E}$ and $G_{D}$ indicates that $\beta=1.882\pm 0.254~{}(13.5\%)$. As intuited, the relation Eq. 18 thus shows that the model described by Eq. 17 only depends in fact on the unique parameter $\xi$. Note that for sake of completeness, the dataset shown in Fig. 6 is supplemented by simulations at various amplitudes, $0.05\leq\gamma_{m}\leq 0.15$, in contrast with Fig. 4 where all simulations are done at $\gamma_{m}=0.15$ only.
@@ 4.4 Loss factor
Here, we provide details about energy dissipation, in line with Sec. 3.4, at the light of the dimensional analysis provided in Sec. 4.3. In particular, we aim at probing how the dissipated energy Eq. 6, the stored energy Eq. 7 and the loss factor Eq. 8 depend on the non-dimensional parameter $\xi$. For this purpose, we compare in Fig. 7 the data simulated by DEM to the features of the reduced model. The mechanical response of the model, $\mathrm{T}(\Gamma)$, is determined via explicit numerical time integration of Eqs. 17-18 for every $\xi$; then one extracts the energetic features according to Eqs. 6-8. Note that for dimensional purpose, energies will be represented with respect to a reference energy,
@@
As expected from the analysis of Sec. 4.3, we observe on Fig. 7 that both energies and the loss factor are well correlated to $\xi$. The agreement between the data simulated by DEM and the reduced model is both qualitative and quantitative in the moderate range of $\xi$. From low-to-moderate $\xi$, the non-dimensional dissipated energy rises because increasing the pressure, for instance, raises the frictional dissipative force $\mu_{D}P$; concurrently, the medium transits from a loose (at low Coulomb threshold) to a dense (at high Coulomb threshold) state, thus progressively gaining elasticity and the ability to store energy. In contrast at higher pressure, the particles firmly stick with each other and the sample tends to store energy as a purely elastic medium: from moderate-to-high $\xi$, the dissipated energy vanishes while the non-dimensional stored energy tends to unity. Consistently with these trends, the loss factor $\eta$ appears relatively constant in the moderate range of $\xi$ (see the colored markers in Fig. 7c, $0.06\leq\xi\leq 0.4$), in agreement with the observation in Fig. 4d that $\eta$ appears fairly independent of the pressure within the same range of parameters. The trends also reveal a decrease of $\eta$ which is characteristic of an elastic medium at high $\xi$, but in a range that is not accessed by the simulations. However, the reduced model reveals a plateau at low $\xi$, but again not evidenced by the DEM simulations.
@@
On the one hand, we supplement six data at high $\xi$ (white squares in Fig. 7) which tend to confirm the decay predicted by the model, namely, a jammed granular medium under high pressure behaves as an elastic solid (weak dissipation, high stored energy and thus low loss factor). On the other hand, the trend of the three data added at small $\xi$ (white triangles in Fig. 7) reveals a decay of $\eta$ which clearly differs from the prediction given by the reduced model. Interestingly, such a decay is concomitant with the nonlinear stiffening inflection observed near the cusp points in Fig. 1e, at low $P$ and high $\gamma_{m}$, concurrently with the significant relative fluctuation of the pressure shown in Fig. 1c. The buildup of such additional nonlinearity in the stress-strain response at low $\xi$ is clearly evidenced in the examples $\xi=0.124$ to $\xi=0.039$ given in the inset of Fig. 7c.
@@ 4.5 Weakly nonlinear response
As expected in the conclusions of Sec. 3.2 and from the developments of Sec. 4.1, the fluctuation of the pressure due to the frustrated dilatancy described in Eq. 3 is a potential source of nonlinearity of the elasticity, and thus an opportunity to improve the reduced model defined in Eqs. 4 and 5. In a first qualitative attempt, aiming at capturing such a nonlinearity in the stress-strain relation, we will only consider a correction in the effective elasticity defined in Eqs. 10-12, by arbitrarily leaving unchanged the pressure in the Coulomb stress, $\mu_{D}P=\mathrm{const}$ see Eq. 5. The first order Taylor expansion of the shear modulus defined in Eq. 13 provides
$G_{T}(\widetilde{P})$ $=$ $G_{T}(P)\times\left(1+\delta P/P\right)^{1/3},$
@@
As an example, we examine the contribution of such a correction for the case $(\phi,\gamma_{m})=(0.64,0.15)$ shown in Fig. 1e, for which $\xi=0.080$ and $\rho=2\Delta P/3P\simeq 0.131$, and whose DEM simulation is reproduced in Fig. 8a. As seen in Fig. 8b, the stress-strain relation predicted by the model defined in Eq. 4, see the red dashed-dotted curve, suffers from a poor agreement: the identification of the parameters described in Sec. 3.3 inherently matches the dissipated energy (see the area within the curve) and well catches an overall elasticity, but fails to reflect neither the slope at small strains (near $\Gamma=0$) nor the characteristic inflections at large strains (near $\Gamma=1$). The Eq. 21 provides a more convincing qualitative agreement, with $G_{NL}=\rho G_{T}/\gamma_{m}^{2}$ estimated by using the value of $G_{T}$ fitted by Eq. 4 and the value $\rho=0.131$ estimated from the pressure fluctuation, see the solid black line in Fig. 8b. However, the quantitative agreement is less obvious: the original model Eq. 4 overestimates the loss factor simulated by DEM by $19$% whereas the improved model Eq. 21 underestimates it by $27$%.
@@
Figure 8: (a) Non-dimensional stress-strain curve of DEM simulation in the weakly nonlinear regime ($\phi=0.64$ at $\gamma_{m}=0.15$, such that $\xi=0.080$ and $\rho=2\Delta P/3P\simeq 0.131$): the DEM simulated loss factor is $\eta=0.368$. (b) Prediction given by the reduced model Eq. 4 (red dashed-dotted line, loss factor $\eta=0.438$) and by the improved reduced model Eq. 21 (black solid line, loss factor $\eta=0.270$). Compared to the DEM simulation, the former does not show the characteristic inflection near the cusp points, underestimates the peak amplitude of the stress by $11$% and overestimates the loss factor by $19$%; the latter qualitatively shows such an inflection, but overestimates the peak amplitude of the stress by $45$% and underestimates the loss factor by $27$%.
@@
A better quantitative agreement can be reached by adjusting the parameters of Eq. 21 independently, since for instance the value of $G_{T}$ obtained with Eq. 4 is poorly relevant in the weakly nonlinear regime. However, the identification of a set of four parameters, $(\mu_{D},G_{D},G_{E},G_{NL})$, introduces numerical uncertainties that require a protocol as careful as the one detailed in Fig. 3. An even more reliable improvement is to consider also the pressure fluctuation not only on the elastic contribution, as in Eq. 21, but in the hysteretic part too. However, both these additional improvements are beyond the scope of our qualitative attempt.
@@
Nevertheless, the dimensional analysis of Eq. 21, derived as in Sec. 4.3, straightforwardly shows that it depends on the two non-dimensional parameters $\xi$ and $\rho$ only. Attempting to reduce further the description of the model, we consider that $\xi$ is essentially a function of $P/k_{n}$ and $\gamma_{m}$ and we probe the correlations of $\rho$ with the same two parameters, as detailed in Figs. A2c and A2d of the Appendix Sec. A2. Understanding such dependencies is out of the scope of the present paper. However, they allow estimating an improved version of the loss factor in a smaller parameter space, $\eta(\xi(P/k_{n},\gamma_{m}),\rho(P/k_{n},\gamma_{m}))$. This improvement is exemplified in Fig. 7c (dashed-dotted black curve) for the nominal case $\gamma_{m}=0.15$ and the pressure dependent $\rho$ as defined on Fig. A2c. The plot shows a trend of the loss factor at low $\xi$ that qualitatively agrees with DEM simulations (see the hollow triangles calculated at low pressures), thus demonstrating that the lowering of $\eta$ likely results from the nonlinearity of the elasticity near the unjamming condition.
@@
From the practical point of view, the decay of the loss factor $\eta$ at both low and high $\xi$ leaves a decade wide region, around $0.50.05<\xi<$ 0.5, where the loss factor is maximal and ranges in between 20% and 40%. Technically, this provides an opportunity to tune the dissipation in order to target specific operating points, by adjusting for instance the confinement pressure to mitigate efficiently a given range of vibrations amplitudes. Interestingly, the optimal value $\eta\simeq 40$% is independent of the particle properties.

### v2

@@ 4.1 Effective elasticity
It is observed in Fig. 4a that the identified elastic moduli $G_{D}$ and $G_{E}$ are both correlated to the confinement pressure $P$. This behavior is expected in granular material. Indeed, focusing on the low amplitude regimes, such that $\tau_{D}\ll\mu_{D}P$, the asymptote of the reduced model defined in Eq. 4 and 5 indicates that $\tau\rightarrow(G_{E}+G_{D})\gamma$}. Hence, a total shear modulus of the granular packing,
@@
is enhanced due to the contribution of the inter-particle contact stiffness in the tangential direction, $k_{t}$, defined in Eq. 2. The expressions Eqs. 10-12 indicate the typical scaling $G_{\mathrm{EMT}}^{0,\infty}\propto K_{\mathrm{EMT}}\propto E_{p}(P/E_{p})^{1/3}$ passed on from Hertzian contact interaction. Finally, it is worth reminding that the volume fraction and the coordination number are also pressure dependent (O’Hern et al., 2003), $\phi(P)$ and $Z(P)$, as pointed out in Sec. 3.5.
@@
with a well-defined coefficient of proportionality, $\alpha=0.687\pm 0.073~(10.6\%)$, on average over a dataset which is restricted to an inter-particle friction coefficient $\mu_{p}=0.15$. In line with the EMT model, the value of the prefactor relies on the inter-particle friction coefficient: the formula given by Eqs. 11-12 indicates two asymptotic values $\alpha=G_{\mathrm{EMT}}^{0}/G_{\mathrm{EMT}}^{\infty}=1/(1+3k_{t}/2k_{n})\simeq 0.4$ for frictionless particles and $\alpha=G_{\mathrm{EMT}}^{\infty}/G_{\mathrm{EMT}}^{\infty}=1$ for frictional ones. The value of $\alpha$ at an intermediate $\mu_{p}=0.15$ is thus consistent with these bounds.
Figure 5: (a): Total shear modulus $G_{T}=G_{D}+G_{E}$ versus effective shear modulus $G_{\mathrm{EMT}}^{\infty}$ (Eq. 12) with a linear fit of slope $\alpha=0.687\pm 0.073~(10.6\%)$. The same legend as in Fig. 4. (b,c): identified Dahl friction coefficient $\mu_{D}$ and coefficient $\alpha$ as functions of the inter-particle friction $\mu_{p}$.
@@ 4.2 Effective friction
which is known to depend on a single dimensionless inertial number (Da Cruz et al., 2005; Jop et al., 2006), $I=t_{p}/t_{s}$. The number $I$ is the ratio between two time scales, namely, the typical time $t_{s}=1/\dot{\gamma}$ associated with shear rate, which correspond to the macroscopic deformation duration, and $t_{p}=r_{p}/\sqrt{P/\rho_{p}}$ that related to confinement, corresponding to the typical duration of microscopic rearrangement during which a grain moves by one neighbor size due the confining pressure (Jop, 2015). The ratio $I$ describes the regime of a granular flow, from the quasi-static limit at low $I$ to the kinetic regime with multiple collisions at high $I$. A phenomenological equation (Jop et al., 2006) provides an estimation of the effective friction coefficient,
@@
where $I_{0}$ is a constant of the order of unity, which delimits the two asymptotic limits. The effective friction coefficient value $\mu_{g}$ accounts for the existence of a gaseous phase at high inertia (Jop, 2015), $I\gg I_{0}$. In turn, the quasi-static value of the effective friction coefficient, $\mu_{s}$, mostly depends on the inter-particle friction coefficient $\mu_{p}$ (Badetti et al., 2018). In practice, the estimated inertial number probed in our study remains small, $10^{-5}\leq I\leq 10^{-4}$, thanks to the low driving frequency (see Sec. 2), in such a way that the effective friction does not depend on the inertial number but only on the inter-particle friction coefficient, $\mu(I\ll I_{0})\simeq\mu_{s}(\mu_{p})$. The analysis of the dependence of Dahl’s effective friction coefficient, $\mu_{D}$ as a function of the inter-particle one, $\mu_{p}$, is shown in Fig. 5b, all the other parameters being fixed. In particular, we observe that $\mu_{D}$ is an increasing function of $\mu_{p}$ with a saturation, in qualitative and quantitative agreement with the literature, see for instance Ref. (Badetti et al., 2018).
Finally, we check the dependence of the prefactor $\alpha$ defined in Eq. 13 with respect to $\mu_{p}$. As expected in Sec. 4.1, $\alpha$ should be a function of $\mu_{p}$, accounting for the transition between perfectly frictionless, $\alpha\simeq 0.4$, and infinitely frictional particles, $\alpha=1$. The observation reported in Fig. 5c is in qualitative agreement with this expectation, showing that the prefactor $\alpha$ increases with the inter-particle friction coefficient $\mu_{p}$ until saturating at a value of the order of unity.
@@ 4.3 Dimensional analysis
We determined in Sec. 3.5 that the mechanical response depends on numerous particle features ($E_{p}$, $\nu_{p}$, $\rho_{p}$, $r_{p}$, $N_{p}$ and $\mu_{p}$), state variables ($P$, $Z$, $\phi$) and external parameters ($\omega_{0}$ and $\gamma_{m}$). However, Sec. 4.1 and 4.2 show that most of these parameters are linked with each other. Here, we aim at reducing the dimension of the parameter space, by making use of the observations presented in Sec. 3.3. In particular, we use the fact that the stress strain response is well described by an elasto-hysteretic model given by Eq. 4 and 5. We thus choose to rewrite these equations in dimensionless form using the following quantities:
@@
with $\Gamma$ and $\mathrm{T}$ the dimensionless shear strain and stress, respectively normalized by the amplitude of the shear strain and a typical shear stress proportional to the total shear modulus $G_{T}$ defined in Eq. 9. In line with this choice, it stands to reason that Dahl’s effective friction coefficient $\mu_{D}$ can be represented by a non-dimensional parameter $\xi$, the first parameter of our model, resulting from the competition between the typical shear stress imposed on the granular medium, $G_{T}\gamma_{m}$, and a Mohr-Coulomb-like threshold $\mu_{D}P$. According to Fig. 4, the parameter $\xi$ is essentially a function of $P/k_{n}$ and $\gamma_{m}$. In turn, the elastic term added to the hysteretic contribution, see Eqs. 4 and 5, can be described by a second parameter, $\Psi$, which stands as a partition parameter quantifying the ratio between the purely elastic shear modulus, $G_{E}$, and that stemming from the hysteretic response, $G_{D}$. According to the normalizations defined in Eq. 16, the model Eqs. 4 and 5 is rewritten as
@@
revealing that the mechanical response $\mathrm{T}(\Gamma)$ depends only on two parameters, $\xi$ and $\Psi$. From a physical point of view, the parameter $\xi$ provides information on the degree of hysteresis (related to frictional stress, $\mu_{D}P$) and elasticity (related to elastic stress, $G_{T}\gamma_{m}$) of the response, suggesting that $\Psi$ should also be a function of $\xi$.
@@
Such an expectation is checked in Fig. 6, which reveals a correlation between $\Psi$ and $\xi$. In practice, the dependency is probed in the small-to-moderate values of $\xi$ only (see Fig. 1d for instance), where a hysteresis is clearly observable and leads to accurate estimations of the triplet $(\mu_{D},G_{D},G_{E})$ i.e. of a pair $(\Psi,\xi)$. For larger $\xi$ (i.e. high pressure $P$ or low amplitude $\gamma_{m}$), the mechanical response tends to an elastic one, which prevents distinguishing the elasto-hysteretic contributions and leads to imprecise estimations of $\Psi$ and $\xi$. However, considering that the granular medium becomes purely elastic in the high $\xi$ regime provides the information that the partition parameter $\Psi$ saturates at 1. Hence, the simplest saturating function takes the form
@@
where $\beta$ is an adjusting parameter. Fitting the Ansatz given in Eq. 18 to both $G_{E}$ and $G_{D}$ indicates that $\beta=1.882\pm 0.254~(13.5\%)$. As intuited, the relation Eq. 18 thus shows that the model described by Eq. 17 depends only on the unique parameter $\xi$. Note that for the sake of completeness, the dataset shown in Fig. 6 is supplemented by simulations at various amplitudes, $0.05\leq\gamma_{m}\leq 0.15$, in contrast with Fig. 4 where all the simulations are done at $\gamma_{m}=0.15$ only.
@@ 4.4 Loss factor
Here, we provide details about energy dissipation, in line with Sec. 3.4, in the light of the dimensional analysis provided in Sec. 4.3. In particular, we aim at probing how the dissipated energy Eq. 6, the stored energy Eq. 7 and the loss factor Eq. 8 depend on the non-dimensional parameter $\xi$. To this end, we compare in Fig. 7 the data simulated by DEM to the features of the reduced model. The mechanical response of the model, $\mathrm{T}(\Gamma)$, is determined via the explicit numerical time integration of Eqs. 17-18 for every $\xi$. Afterwards, the energetic features are extracted according to Eqs. 6-8. Note that for dimensional purposes, energies will be represented with respect to a reference energy,
@@
As expected from the analysis of Sec. 4.3, we observe in Fig. 7 that both energies and the loss factor are well correlated to $\xi$. The agreement between the data simulated by DEM and the reduced model is both qualitative and quantitative in the moderate range of $\xi$. From low-to-moderate $\xi$, the non-dimensional dissipated energy rises because increasing the pressure, for instance, raises the frictional dissipative force $\mu_{D}P$; concurrently, the medium transits from a loose (at low Coulomb threshold) to a dense (at high Coulomb threshold) state, thus progressively gaining elasticity and the ability to store energy. In contrast, at higher pressure the particles stick firmly to each other and the sample tends to store energy as a purely elastic medium: from moderate-to-high $\xi$, the dissipated energy vanishes while the non-dimensional stored energy tends to unity. Consistently with these trends, the loss factor $\eta$ appears relatively constant in the moderate range of $\xi$ (see the colored markers in Fig. 7c, $0.06\leq\xi\leq 0.4$), in agreement with the observation in Fig. 4d that $\eta$ appears fairly independent of the pressure within the same range of parameters. The trends also reveal a decrease of $\eta$ which is characteristic of an elastic medium at high $\xi$, but in a range that is not accessed by the simulations. However, the reduced model reveals a plateau at low $\xi$, but again not evidenced by the DEM simulations.
@@
On the one hand, we supplement six data at high $\xi$ (white squares in Fig. 7) which tend to confirm the decay predicted by the model, namely, a jammed granular medium under high pressure behaves as an elastic solid (weak dissipation, high stored energy, and thus low loss factor). On the other hand, the trend of the three data added at low $\xi$ (white triangles in Fig. 7) reveals a decay of $\eta$ which clearly differs from the prediction given by the reduced model. Interestingly, such a decay is concomitant with the nonlinear stiffening inflection observed near the cusp points in Fig. 1e, at low $P$ and high $\gamma_{m}$, concurrently with the significant relative fluctuation of the pressure shown in Fig. 1c. The buildup of such additional nonlinearity in the stress-strain response at low $\xi$ is clearly evidenced in the examples $\xi=0.124$ to $\xi=0.039$ given in the inset of Fig. 7c.
@@ 4.5 Weakly nonlinear response
As expected in the conclusions of Sec. 3.2 and from the developments of Sec. 4.1, the fluctuation of the pressure due to the restrained dilatancy described in Eq. 3 is a potential source of nonlinearity of the elasticity, and thus an opportunity to improve the reduced model defined in Eqs. 4 and 5. In the first qualitative attempt, aimed at capturing such a nonlinearity in the stress-strain relation, we will only consider a correction in the effective elasticity defined in Eqs. 10-12, by arbitrarily leaving unchanged the pressure in the Coulomb stress, $\mu_{D}P=\mathrm{const}$ see Eq. 5. The first order Taylor expansion of the shear modulus defined in Eq. 13 provides
[…]
@@
As an example, we examine the contribution of such a correction for the case $(\phi,\gamma_{m})=(0.64,0.15)$ shown in Fig. 1e, for which $\xi=0.080$ and $\rho=2\Delta P/3P\simeq 0.131$, and whose DEM simulation is reproduced in Fig. 8a. As seen in Fig. 8b, the stress-strain relation predicted by the model defined in Eq. 4, see the red dashed-dotted curve, suffers from poor agreement: the identification of the parameters described in Sec. 3.3 inherently matches the dissipated energy (see the area within the curve) and acquires overall elasticity well, but fails to reflect either the slope at small strains (near $\Gamma=0$) or the characteristic inflections at large strains (near $\Gamma=1$). The Eq. 21 provides a more convincing qualitative agreement, with $G_{NL}=\rho G_{T}/\gamma_{m}^{2}$ estimated by using the value of $G_{T}$ fitted by Eq. 4 and the value $\rho=0.131$ estimated from the pressure fluctuation, see the solid black line in Fig. 8b. However, the quantitative agreement is less obvious: the original model Eq. 4 overestimates the loss factor simulated by DEM by 1919% whereas the improved model Eq. 21 underestimates it by 2727%.
@@
Figure 8: (a) Non-dimensional stress-strain curve of DEM simulation in the weakly nonlinear regime ($\phi=0.64$ at $\gamma_{m}=0.15$, such that $\xi=0.080$ and $\rho=2\Delta P/3P\simeq 0.131$): the DEM simulated loss factor is $\eta=0.368$. (b) Prediction given by the reduced model Eq. 4 (red dashed-dotted line, loss factor $\eta=0.438$) and by the improved reduced model Eq. 21 (black solid line, loss factor $\eta=0.270$). Compared to the DEM simulation, the former does not show the characteristic inflection near the cusp points, underestimates the peak amplitude of the stress by 1111% and overestimates the loss factor by 1919%. The latter shows this inflection qualitatively, but overestimates the peak amplitude of the stress by 4545% and underestimates the loss factor by 2727%.
@@
A better quantitative agreement can be reached by adjusting the parameters of Eq. 21 independently, since for instance the value of $G_{T}$ obtained with Eq. 4 is not very relevant in the weakly nonlinear regime. However, the identification of a set of four parameters, $(\mu_{D},G_{D},G_{E},G_{NL})$, introduces numerical uncertainties that require a protocol as careful as that detailed in Fig. 3. An even more reliable improvement is to also consider the pressure fluctuation not only on the elastic contribution, as in Eq. 21, but in the hysteretic part too. However, both these additional improvements are beyond the scope of our qualitative attempt.
@@
Nevertheless, the dimensional analysis of Eq. 21, derived as in Sec. 4.3, shows straightforwardly that it depends on the two non-dimensional parameters $\xi$ and $\rho$ only. Attempting to reduce further the description of the model, we consider that $\xi$ is essentially a function of $P/k_{n}$ and $\gamma_{m}$ and we probe the correlations of $\rho$ with the same two parameters, as detailed in Figs. A2c and A2d of the Appendix Sec. A2. Understanding such dependencies is out of the scope of the present paper. However, they allow estimating an improved version of the loss factor in a smaller parameter space, $\eta(\xi(P/k_{n},\gamma_{m}),\rho(P/k_{n},\gamma_{m}))$. This improvement is exemplified in Fig. 7c (dashed-dotted black curve) for the nominal case $\gamma_{m}=0.15$ and the pressure dependent $\rho$ as defined in Fig. A2c. The plot shows a trend of the loss factor at low $\xi$ that agrees with DEM simulations qualitatively (see the hollow triangles calculated at low pressures), thus demonstrating that the lowering of $\eta$ likely results from the nonlinearity of the elasticity near the unjamming condition.
@@
From the practical point of view, the decay of the loss factor $\eta$ at both low and high $\xi$ leaves a decade wide region, around $0.05<\xi<0.5$, where the loss factor is maximal and ranges between 20% and 40%. Technically, this provides an opportunity to tune the dissipation in order to target specific operating points, by adjusting for instance the confinement pressure to mitigate a given range of vibrations amplitudes efficiently. Interestingly, the optimal value $\eta\simeq 40$% is independent of the particle properties.